OpenStudy (anonymous):

Simplify the expression a+1+a-2+a+3+a-4.

OpenStudy (cj49):

4a-2

OpenStudy (cj49):

4a=2 a=1/2

OpenStudy (anonymous):

how did you do it

OpenStudy (cj49):

a+1+a-2+a+3+a-4=0 4a-2=0 4a=0+2 4a=2 a=2/4 a=1/2

OpenStudy (anonymous):

what about this question Simplify $2w+4w+6w+8w+10w+12. help me understand

OpenStudy (cj49):

30w+12=0(divide by 2) 15w+6=0 15w=-6 w=-6/15 w=-2/5

OpenStudy (anonymous):

it says it's incorrect

OpenStudy (cj49):

this is the question ryt Simplify 2w+4w+6w+8w+10w+12.

OpenStudy (anonymous):

that's the question

OpenStudy (cj49):

2w+4w+6w+8w+10W=30w 30w+12=0 30w=-12 w=-12/30 w=-2/5

OpenStudy (anonymous):

so the answer is?

OpenStudy (cj49):

w= -2/5

OpenStudy (anonymous):

oh ok I got it thanks

OpenStudy (cj49):

lol..np..